OpenStudy (anonymous):

how long is a guy wire reaching from the top of a three foot pole to a point on the ground 2 ft from the pole what is the exact length

OpenStudy (anonymous):

the wire, the pole and the ground form a right triangle. Use \[a^{2}+b^{2}=c^{2}\] to solve, where c = the hypotenuse = length of the wire

OpenStudy (anonymous):

\[c^2=2^2+3^2\]

OpenStudy (anonymous):

\[c=\sqrt{13}\]

OpenStudy (anonymous):

i am sorry but you guys are really confusing me, please work these out so i can see how they are done if you could

OpenStudy (anonymous):

|\ | \ | \c 3ft \ | \ | \ |-----2ft--\ 3ft = a , 2ft =b , a^2+b^2=c^2

OpenStudy (anonymous):

what would the length be approximated to three decimal places

OpenStudy (anonymous):

u calculator \[\sqrt{13} \approx 3.605551275\] u round

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i understood what you said better thanks

OpenStudy (anonymous):

urwc